40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 20 12:12:25 2012, rana, Update, PSL, PMC problems examined Untitled.png
    Reply  Sat Jul 21 14:05:21 2012, Den, Update, PSL, PMC problems examined fss_slowm.png
Message ID: 6998     Entry time: Sat Jul 21 14:05:21 2012     In reply to: 6995
Author: Den 
Type: Update 
Category: PSL 
Subject: PMC problems examined 

Quote:

WE found that the PMC EPICS values had not been toggled since the reboot and so the RF phase and Amplitude were totally wrong (we should replace this with a fixed oscillator box as we did with FSS).

Also, the NPRO SLOW slider was at -2 V which made the mode going into the PMC funny (although the mode was OK this morning before I started playing with the PMC sliders).

 PMC transmission is oscillating  in the range 0.5 - 0.85. PMC PZT voltage is 1-2 V.

FSS slow controls was -2.5 V. I adjusted it to 0 and PMC stabilized. PMC PZT voltage is 128, transmission is 0.845.

But most probably, slow control will drift again.

fss_slowm.png

ELOG V3.1.3-